The equation of discharge of a capacitor is given by Q = Qoe^(-t/RC) , define the terms half life and time constant of a discharging capacitor and derive the expression that links them.
1
Expert's answer
2021-02-02T11:48:49-0500
  • The half-time "T" is the time during which the charge stored is decreased by two times : "Q_0 e^{-\frac{T}{RC}} = \frac{1}{2} Q_0". Thus "\frac{T}{RC}=\ln 2" and so "T = RC\ln2".
  • The characteristic time "\tau" is defined such that "Q = Q_0 e^{-\frac{t}{\tau}}" and thus "\tau = RC".

We see from these two relations that "T = \tau \ln 2". We can also verify that "RC" have the dimensions of a time.